OpenStudy (anonymous):

what is the answer to 9x-y=-2

OpenStudy (precal):

just pick an x value how about x=0 so then solve for y

OpenStudy (solomonzelman):

What exactly are you looking for ?

OpenStudy (anonymous):

you can use substition: solve for one variable and sub that value into equation -y=-2-9x y=2+9x now sub this value in place of the y in original equation

OpenStudy (anonymous):

substitution